Sodium sulfite is a common preservative that can be found in wine, dry fruits and vegetables, juices, sausages,and other foods. A new food factory plans to buy 150 gallon tank to storesodium sulfite and has hired you as a consultant to help them determine the maximum mass of sodium sulfite that they can fit inside the tank. Given that the density of sodium sulfite is 2.63 g/mL, how much would 150gallons of it weigh? (1 gal= 4 qt = 3.7854 L,1 lb = 453.59 g)
Answer #1

One gallon is equal to the 3.754 L = 3785.4 mL

We know that density is equal to mass/volume.

Density of sodium sulfite = 2.63 g/ml

Volume of tank = 150 gallons = 150*3.7853 = 567.81 L = 567810 mL

Mass of 150 gallons of sodium sulfite= density * volume

= 2.63*567810=1493340.3 grams.